OpenStudy (anonymous):

Anyone stuck on h8p1?

OpenStudy (anonymous):

We have a circuit with a resistor and inductor in series. a) we need to find I(0-): we use the equation V0/R+{i(-1)-(V0/R]*E^(-T*(R/L)) They give you V0, i(-1), R, L, t=0.001 b)I(0+)=I(0-) + lamda/L c)I(1ms)= V0/R*(1-E^(-T*(R/L))+I(0+)(E^(-T*(R/L)) hope that helps

OpenStudy (anonymous):

v_R1_0min = 0 v_r1_0plus = 2000 v_r1_1ms = 270.6706 v_c_0plus = 3.0303e+003 v_c_0plus = 146.3667